If 80cm cube of methane (CH4)diffuses through a porous membrane in 20 seconds while 50cm cube of gas Q diffuse through the same membrane in 25 seconds.Calculate the molecular mass of gas Q


1
Expert's answer
2022-05-05T02:42:03-0400

R1/R2=√M2/M1. 25*80/50=40




M2=(R1/R2)^2 x M1 = (20/40) x 16 = 4.




He.